# The cosmic-ray sea explains the Galactic -ray and diffuse emissions from GeV to PeV

### v1

@@ Abstract
The LHAASO collaboration has recently released the spectrum and the angular distribution of the gamma-ray Galactic diffuse emission from 1 TeV to 1 PeV measured with the Kilometer-2 Array (KM2A) and the Water Cherenkov Detector Array (WCDA). We show that those data are in remarkably good agreement with a set of pre-existing models that assume the emission to be produced by the Galactic population of cosmic rays if its spectral shape traces that measured by CALET, DAMPE as well as KASCADE at higher energies. No extra-components besides the CR sea is needed to explain LHAASO results. Spatial dependent CR transport models, although not required to reproduce LHAASO results, are in better agreement with them respect to conventional ones and needed to consistently reproduce Fermi-LAT and neutrino data.
@@ 1 Introduction
The Large High Altitude Air Shower Observatory (LHAASO) is a large area, wide field-of-view observatory for CRs and -rays [1]. This experiment combines measurements taken with the Kilometer-2 Array (KM2A) charged particle detector with those of the Water Cherenkov Detector Array (WCDA) allowing to cover an energy interval going from few TeV up to PeV. Due to its wide field of view ( sr), energy band coverage and good angular resolution, LHAASO is presently the best instruments suited to study the -ray Galactic Diffuse Emission (GDE) above the TeV. In 2023 the LHAASO collaboration presented its official results for the average diffuse flux in two different regions covering the Galactic Plane (GP) – the inner () and outer () regions (both with ) – from TeV to PeV obtained with the KM2A array, in Ref. [2]. More recently the collaboration updated those results including also the -GDE flux measurement performed with the WCDA from to TeV, allowing the release of the wide-band spectral distributions in the inner and outer regions from 1 TeV to 1 PeV [4] hence almost connecting to Fermi-LAT results [3]. Moreover LHAASO measured the latitude and longitude GDE profiles in the - TeV and - TeV energy bands. The detection significance measured by the combined WCDAKM2A was found to be 24.6 and 9.1 in the inner and outer regions respectively.
@@
It is important to point-out that in order to suppress the contamination of resolved (especially extended ones) and unresolved sources, LHAASO used a mask which cuts away a large fraction the low Galactic latitude () region of the inner GP. This is different from the approach followed by the Tibet-AS [5] collaboration – which also measured the GDE though only between 100 TeV and 1 PeV and with a smaller significance – which masked only a small area extending around the nominal position of each known point like source. We notice that the GDE found by Tibet-AS is significantly higher than LHAASO, the two measurements being in better agreement only if the latter were determined using no mask [6]. Even taking into account the possible enhancement and hardening of the GDE in the inner GP (see below), this discrepancy suggests that Tibet signal may be contaminated by sources, especially extended ones (see also [2]).
@@
LHAASO results represent an important achievement for astroparticle physics. In fact, the GDE is a background for point like and, more critically, extended source observations as well as for dark matter indirect search. More importantly in the present context, its measurement provides an unique handle to determine the spectrum and the large scale spatial distribution of the Galactic Cosmic Ray (CR) population while direct detection of charged particles, with satellite and ground detectors, only probe relatively local CR properties. Moreover, at energies above few hundred TeV these properties can only be measured indirectly studying the secondary particle showers they produce in the atmosphere. That procedure introduces large systematics mostly connected to the simulation of the air showers development. In this regard we observe the large offset – still compatible with systematics (see Ref. [7]) – of the CR proton spectrum measured by KASCADE [8] respect to that published by IceTop [9] and, very recently, by the LHAASO collaboration [10]. Waiting for experimental groups to get to a better agreement, the newly born PeV -ray astronomy offers an independent insight especially if exploited in combination with high energy neutrino astronomy so to get rid of the degeneracy between leptonic and hadronic -ray emissions.
@@
We stress, however, that even precisely knowing the CR spectra at the Earth position, is not sufficient to allow a realistic modelling of the GDE which is produced by the interaction of the CR hadrons and electrons with the interstellar gas and radiation field respectively. In fact, due to not uniform spatial distribution of CR sources and to the shape of the Galaxy, also the CR large scale distribution is not expected to be uniform but rather peaked around the inner molecular ring, a couple of kpc away from the Galactic Centre (GC). Furthermore, observational [11, 3] and theoretical arguments [12] suggest that CR transport may also be spatial dependent giving rise to a hardening of the CR spectrum towards the GC. Accounting for the – also quite peaked towards the GC – gas distribution all this can have a major impact on the GDE at energies above the TeV.
In order to account for these, as well as other, effects so to produce as much as possible realistic and accurate emission maps to be used as templates by the current high resolution experiments, dedicated numerical codes have to be used. Here we use the DRAGON code [13] which in combination with HERMES [14] allows to model the and GDEs under the condition it reproduces the primary and secondary CR spectra measured at the Earth. High energy -ray attenuation due to the scatteting on the CMB is properly taken into account.
@@ 2 Conventional versus space dependent CR transport
First of all we recall that the LHAASO-WCDA measured different GDE spectra in the inner () and outer () regions finding them to follow power-laws with indexes and respectively [4]. This provides a mild hint (with a significance of ) that the primary CR spectrum may get harder toward the GC. That however is not enough to reach a firm conclusion.
In Fig. 3 we compare the Base (our reference conventional model) and -optimized models – using for both the Min source spectrum setup – with the LHAASO-WCDA data. Note that for consistency the LHAASO mask is applied to the models as well as to the Femi-LAT data explaining the degeneracy at low energy (see below). Again, the LHAASO data looks in better agreement with the spatial dependent models although – also due to the absence of the, yet not released, systematic errors – the significance of that preference is still too low to draw robust conclusions.
While one would aspect the offset between conventional and space-dependent models to grow even more with energy, it should taken in mind that LHAASO masks a large fraction the inner GP region which is where the predictions of the two scenarios are most significant. That effect is clearly displayed in Fig.2. As a consequence, we notice that the Fig.1 would not change significantly replacing the curve models with those obtained for the corresponding Base models.
@@ 3 Implications for neutrino astronomy
Interestingly, the IceCube collaboration found its best-fit to be in good agreement with the -GDE measured by Tibet-AS [5]. Unfortunately, as discussed in the above, the GDE spectrum measured by Tibet is not compatible with the – more reliable – LHAASO results and it is likely to be contaminated by source emissions. This is especially true taking in mind that conventional models do not predict a large enhancement of the GDE in the GP region masked by LHAASO. Last but not least the model is disfavoured by IceCube results themselves. In fact, the model independent pre-trial significance map of the all-sky search and the best-fit spectral index angular distribution show the emission to be more concentrated and harder than predicted by that model (see the Supplementary material in [19]).
@@ 4 Conclusions
We showed that the spectrum and the angular distribution of the -ray diffuse emission of the Galaxy recently measured by the LHAASO collaboration are naturally interpreted in terms of secondary emission by the Galactic CR population if their spectral shape traces that measured by KASCADE (Min setup).
As long as the mask adopted by LHAASO is used also for Fermi-LAT, both data sets are reproduced by conventional (Base) as well as spatial-dependent (-optimized) CR transport models. We have shown the degeneracy to be a consequence of that mask cutting the inner GP region where the predictions of those models mostly differ. Hopefully, in the next future LHAASO may refine its masking procedure so to better cover the inner Galactic plane region breaking that degeneracy. The SWGO water Cherenkov experiment [24], which will provide a more direct view of the GC region, will provide a valuable complementary scrutiny. It should be stressed, however, that only the -optimized models allow to reproduce Fermi-LAT results and should therefore be preferred providing a comprehensive description of the Galactic diffuse emission. That scenario seems also to be favoured by IceCube neutrino measurements which found a hint of a spectral hardening in the -GDE emission [19]. While IceCube found already a good fit using the spatial-dependent KRA model as a template, we argued that agreement may improve using its -optimized Min update.

### v2

@@ Abstract
The LHAASO collaboration has recently released the spectrum and the angular distribution of the gamma-ray Galactic diffuse emission from 1 TeV to 1 PeV measured with the Kilometer-2 Array (KM2A) and the Water Cherenkov Detector Array (WCDA). We show that those data are in remarkably good agreement with a set of pre-existing models that assume the emission to be produced by the Galactic population of cosmic rays if its spectral shape traces that measured by CALET, DAMPE as well as KASCADE at higher energies. No extra-component besides the CR sea is needed to explain LHAASO results. Spatial dependent CR transport models, although not required to reproduce LHAASO results, are in better agreement with them respect to conventional ones and needed to consistently reproduce Fermi-LAT and neutrino data.
@@ 1 Introduction
The Large High Altitude Air Shower Observatory (LHAASO) is a large area, wide field-of-view observatory for CRs and -rays [1]. This experiment combines measurements taken by the Kilometer-2 Array (KM2A) charged particle detector with those of the Water Cherenkov Detector Array (WCDA) allowing to cover an energy interval going from few TeV up to PeV. LHAASO is presently the best instrument suited to study the -ray Galactic Diffuse Emission (GDE) above the TeV. In 2023 the LHAASO collaboration released its measurement of the average diffuse flux in two different regions covering the Galactic Plane (GP) – the inner () and outer () regions (both with ) – from TeV to PeV obtained with the KM2A array, in Ref. [2]. More recently the collaboration updated those results including also the -GDE flux measurement performed with the WCDA from to TeV, allowing the release of the wide-band spectral distribution from 1 TeV to 1 PeV [4] hence almost connecting to Fermi-LAT results [3]. Moreover LHAASO measured the latitude and longitude GDE profiles in the - TeV and - TeV energy bands. The detection significance measured by the combined WCDAKM2A was found to be 24.6 and 9.1 in the inner and outer regions respectively.
@@
It is important to point-out that in order to suppress the contamination of resolved (especially extended ones) and unresolved sources, LHAASO used a mask which cuts away a large fraction the low Galactic latitude () region of the inner GP. This is different from the approach followed by the Tibet-AS [5] collaboration – which also measured the GDE though only between 100 TeV and 1 PeV and with a smaller significance – which masked only a small area extending around the nominal position of each known point like source. We notice that the GDE found by Tibet-AS is significantly higher than LHAASO, the two measurements being in better agreement only if the latter were determined using no mask [6]. Even taking into account the possible enhancement and hardening of the GDE in the inner GP (see below), this discrepancy suggests that Tibet signal may be partially contaminated by sources, especially extended ones (see also [2]).
@@
LHAASO results represent an important achievement for astroparticle physics. In fact, the GDE is a background for point like and, more critically, extended source observations as well as for dark matter indirect search. More importantly in the present context, its measurement provides an unique handle to determine the spectrum and the large scale spatial distribution of the Galactic Cosmic Ray (CR) population while direct detection of charged particles, with satellite and ground detectors, only probe relatively local CR properties. Moreover, at energies above few hundred TeV these properties can only be measured indirectly studying the secondary particle showers they produce in the atmosphere. That procedure introduces large systematics mostly connected to the simulation of the air showers development. In this regard we observe the large offset of the CR proton spectrum measured by KASCADE [7] respect to that published by IceTop [8] and, very recently, by the LHAASO collaboration [9]. Waiting for experimental groups to get to a better agreement, the newly born PeV -ray astronomy offers an independent insight especially if exploited in combination with high energy neutrino astronomy so to get rid of the degeneracy between leptonic and hadronic -ray emissions.
@@
We stress, however, that even precisely knowing the CR spectra at the Earth position, is not sufficient to allow a realistic modelling of the GDE which is produced by the interaction of the CR hadrons and electrons with the interstellar gas and radiation field respectively. In fact, due to not uniform spatial distribution of CR sources and to the shape of the Galaxy, also the CR large scale distribution is not expected to be uniform but rather peaked around the inner molecular ring, a couple of kpc away from the Galactic Centre (GC). Furthermore, observational [10, 3] and theoretical arguments [11] suggest that CR transport may also be spatial dependent giving rise to a hardening of the CR spectrum towards the GC. Accounting for the – also quite peaked towards the GC – gas distribution all this can have a major impact on the GDE at energies above the TeV. In order to account for these, as well as other, effects so to produce as much as possible realistic and accurate emission maps to be used as templates by the current high resolution experiments, dedicated numerical codes have to be used. Here we use the DRAGON code [12] which in combination with HERMES [13] allows to model the and GDEs under the condition it reproduces – as a working hypothesis to be tested at posteriori – the primary and secondary CR spectra measured at the Earth. High energy -ray attenuation due to the scatteting on the CMB is properly taken into account.
@@ 2 Conventional versus space dependent CR transport
First of all we recall that the LHAASO-WCDA measured different GDE spectra in the inner () and outer () regions finding them to follow power-laws with indexes and respectively [4]. This provides a mild hint (with a significance of ) that the primary CR spectrum may get harder toward the GC.
In Fig. 3 we compare our models with the masked LHAASO-WCDA data. Again, the LHAASO data looks in better agreement with the spatial dependent models although – also due to the absence of the, yet not released, systematic errors – the significance of that preference is still too low to draw robust conclusions.
While one would aspect the offset between conventional and space-dependent models to grow even more with energy, it should taken in mind that LHAASO masks a large fraction the inner GP region which is where the predictions of the two scenarios are most significant. That effect is clearly displayed in the left panels of Fig.2. A stronger discriminating power is provided by unmasked model and LHAASO data (right panels). That, however, is meaningful only above 100 TeV where the contamination of unresolved sources is expected to be suppressed. Interestingly, a good agreement – improving as the energy approaches the PeV – is found between those data and the -optimized Min model.
@@ 3 Implications for neutrino astronomy
While the IceCube collaboration found its best-fit to be in good agreement with the -GDE measured by Tibet-AS [5], unfortunately, as discussed in the above, the GDE spectrum measured by Tibet is not compatible with the LHAASO results and it is likely to be contaminated by source emissions. This is especially true taking in mind that conventional models do not predict a large enhancement of the GDE in the GP region masked by LHAASO. Last but not least the model is disfavoured by IceCube results themselves. In fact, the model independent pre-trial significance map of the all-sky search and the best-fit spectral index angular distribution show the emission to be more concentrated and harder than predicted by that model (see the Supplementary material in [18]).
@@ 4 Conclusions
We showed that the spectrum and the angular distribution of the -ray diffuse emission of the Galaxy recently measured by the LHAASO collaboration are naturally interpreted in terms of secondary emission by the Galactic CR population if their spectral shape traces that measured by KASCADE (Min setup). As long as the mask adopted by LHAASO is used also for Fermi-LAT, both data sets are reproduced by conventional (Base) as well as spatial-dependent (-optimized) CR transport models. We have shown the degeneracy to be a consequence of that mask cutting the inner GP region where the predictions of those models mostly differ. Hopefully, in the next future LHAASO may refine its masking procedure so to better cover the inner Galactic plane region removing that degeneracy. Interestingly, we found that unmasked LHAASO data already favour the -optimized Min model. Moreover, it should be stressed that the -optimized models are in better agreement with Fermi-LAT results and should therefore be preferred providing a comprehensive description of the Galactic diffuse emission. The SWGO water Cherenkov experiment [23], which will allow a closer view of the GC region, may provide a valuable complementary scrutiny. The space-dependent transport scenario seems also to be favoured by IceCube neutrino measurements which found a hint of a spectral hardening in the -GDE emission [18]. While IceCube already found a very good fit () using the spatial-dependent KRA model as a template, we argued that agreement may further improve using its -optimized Min update.